# Mechanically Interlocked Polymers in Dilute Solution under Shear and Extensional Flows: A Brownian Dynamics Study

### v1

@@ V Conclusions
In this study, we explored how mechanical bonds affect the dynamics and rheology of dilute solutions of mechanically-interlocked polymers under steady shear and uniaxial extension. We used a coarse-grained bead-spring model with excluded volume and hydrodynamic interactions in a Brownian dynamics simulation. To elucidate the effect of mechanical interlocks on properties, we ran a series of simulations across a range of flow strengths (Wi numbers) and polymer structures. In polyrotaxanes, the number of rings was varied. In daisy chains, the number and the length of the linear segments were varied. Lastly, in polycatenanes, the number of rings and their sizes were varied. Our simulations reveal that mechanical interlocks significantly alter polymer behavior under various flow conditions, with polyrotaxanes, daisy chains, and polycatenanes each exhibiting distinctive rheological signatures compared to linear polymers. This work provides a foundation for understanding MIP behavior in dilute solutions under weak to moderate flows. Addressing strong flow regimes and mechanical degradation mechanisms in future studies will help fully realize the potential of these novel materials.
acknowledgments
@@
Supplementary Material
The supplementary material includes the fractional extension in $y$ and $z$ directions, orientation angle, and first normal stress coefficient against Wi; equilibrium properties of the polymers studied; comparison graphs of linear and MIPs with similar contour length.
Author Declarations
Conflict of Interest
The authors have no conflicts.
Data Availability
@@ Appendix A Equilibrium Properties
     $M_{w}$      $N_{bb}$      $N_{\mathrm{rings}}$      $R_{g}^{2}/a^{2}$      $D_{COM}\zeta/k_{B}T$      $\tau_{R}/\tau$
     56      40      2      119.98      0.1433      139.57
     72      40      4      125.82      0.1373      152.78
     88      40      6      133.52      0.1312      169.60
     104      40      8      143.14      0.1256      189.94
     112      80      4      237.04      0.0992      398.37
     144      80      8      263.23      0.0932      470.75
     176      80      12      303.02      0.0871      579.55
     208      80      16      341.24      0.0823      690.90
     168      120      6      349.41      0.0801      727.28
     216      120      12      455.36      0.0730      1040.23
     264      120      18      507.79      0.0686      1234.32
     312      120      24      567.62      0.0643      1470.45
@@
     $M_{w}$      BPS      $N_{\mathrm{seg}}$      $R_{g}^{2}/a^{2}$      $D_{COM}\zeta/k_{B}T$      $\tau_{R}/\tau$
     30      8      2      20.49      0.2504      13.64
     74      8      4      61.01      0.1624      62.61
     118      8      6      116.86      0.1254      155.26
     206      8      10      247.67      0.0914      451.62
     46      16      2      46.10      0.1874      41.01
     106      16      4      125.10      0.1228      169.84
     166      16      6      202.98      0.0972      347.90
     226      16      8      307.33      0.0811      631.57
     78      32      2      103.86      0.1331      130.08
     170      32      4      279.14      0.0872      533.57
     262      32      6      472.17      0.0687      1145.30
@@
     $M_{w}$      BPR      $N_{\mathrm{rings}}$      $R_{g}^{2}/a^{2}$      $D_{COM}\zeta/k_{B}T$      $\tau_{R}/\tau$
     16      8      2      7.12      0.3544      3.35
     40      8      5      21.49      0.2396      14.91
     80      8      10      54.14      0.1685      53.74
     168      8      21      145.80      0.1122      217.18
     344      8      43      373.70      0.0745      835.88
     520      8      65      621.31      0.0590      1753.85
     160      16      10      117.34      0.1133      172.67
     240      24      10      184.25      0.0898      341.97
     320      32      10      275.50      0.0751      611.21
@@ Appendix B Additional Shear Flow Properties
Figure B.1: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, (c) orientation angle ($\theta$), and (d) first normal stress coefficient ($\Psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
@@
Figure B.2: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, (c) orientation angle ($\theta$), and (d) first normal stress coefficient ($\Psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate more segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
@@
Figure B.3: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (f) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, (c) orientation angle ($\theta$), and (d) first normal stress coefficient ($\Psi_{1}$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$).
@@
Figure B.4: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, (c) orientation angle ($\theta$), and (d) first normal stress coefficient ($\Psi_{1}$). Darker shades indicate higher beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.
@@
Figure B.5: Comparison of linear and MIP polymers of similar maximum contour length (L) in extensional (a, b) and shear (c, d) flow. Fractional extension in the flow direction $(\langle X_{bb}\rangle/L)$ for extension (a) and shear (c) along with extensional viscosity $(\eta_{e})$ (b) and shear viscosity $(\eta)$ are given as a function of Weissenberg number (Wi). The linear polymer (black) has 80 backbone beads. The polyrotaxane (blue) has 80 backbone beads including 2 capping beads and 16 rings. The daisy chain (green) has 6 linear segments of 8 beads each interlocked with 10 8 bead rings and includes 2 capping beads. The polycatenane (red) has 10 interconnected rings of 16 beads each.

### v2

@@ V Conclusions
In this study, we explored how mechanical bonds affect the dynamics and rheology of dilute solutions of mechanically-interlocked polymers under steady uniaxial extension and shear. We used a coarse-grained bead-spring model with excluded volume and hydrodynamic interactions in a Brownian dynamics simulation. To elucidate the effect of mechanical interlocks on properties, we ran a series of simulations across a range of flow strengths (Wi numbers) and polymer structures. In polyrotaxanes, the number of rings was varied. In daisy chains, the number and the length of the linear segments were varied. In polycatenanes, the number of rings and their sizes were varied. Our simulations reveal that mechanical interlocks significantly alter polymer behavior under various flow conditions, with polyrotaxanes, daisy chains, and polycatenanes each exhibiting distinctive rheological signatures compared to linear polymers. In particular we note that MIPs generally have greater tumbling rates in shear flow and weaker shear thinning compared to linear polymers, while differences in extension depend on flow strength, type, and polymer architecture. This work provides a foundation for understanding MIP behavior in dilute solutions under weak to moderate flows. Addressing strong flow regimes and mechanical degradation mechanisms in future studies will help fully realize the potential of these novel materials.
VI acknowledgments
@@
VII Supplementary Information
The supplementary information includes linear and MIP equilibrium properties (diffusion coefficient, relaxation time, radius of gyration, asphericity, prolateness); additional shear flow measurements as a function of Wi (gradient and cross-flow extent, orientation angle); comparisons of linear and MIPs with similar contour length; trajectory snapshots; probability distribution of rings in polyrotaxanes; and power-law exponents ($\beta$) of $\eta$ and $\psi_{1}$ fitted from shear data.
VIII Data Availability
@@ Appendix A Equilibrium Properties
$M_{w}$ $N_{bb}$ $N_{\mathrm{rings}}$ $D_{COM}\zeta/k_{B}T$ $\tau_{R}/\tau$ $R_{g}/a$ Asphericity Prolateness
56 40 2 0.1433 139.57 10.95 0.73 0.82
72 40 4 0.1373 152.78 11.22 0.72 0.80
88 40 6 0.1312 169.60 11.56 0.73 0.84
104 40 8 0.1256 189.94 11.96 0.75 0.86
112 80 4 0.0992 398.37 15.40 0.64 0.57
144 80 8 0.0932 470.75 16.22 0.63 0.55
176 80 12 0.0871 579.55 17.41 0.66 0.63
208 80 16 0.0823 690.90 18.47 0.66 0.62
168 120 6 0.0801 727.28 18.69 0.61 0.49
216 120 12 0.0730 1040.23 21.34 0.67 0.65
264 120 18 0.0686 1234.32 22.53 0.64 0.59
312 120 24 0.0643 1470.45 23.80 0.62 0.53
@@
$M_{w}$ BPS $N_{\mathrm{seg}}$ $D_{COM}\zeta/k_{B}T$ $\tau_{R}/\tau$ $R_{g}/a$ Asphericity Prolateness
30 8 2 0.2504 13.64 4.53 0.61 0.47
74 8 4 0.1624 62.61 7.81 0.70 0.74
118 8 6 0.1254 155.26 10.79 0.71 0.76
206 8 10 0.0914 451.62 15.75 0.68 0.70
46 16 2 0.1874 41.01 6.79 0.62 0.52
106 16 4 0.1228 169.84 11.18 0.66 0.63
166 16 6 0.0972 347.90 14.26 0.62 0.53
226 16 8 0.0811 631.57 17.57 0.63 0.55
78 32 2 0.1331 130.08 10.19 0.61 0.49
170 32 4 0.0872 533.57 16.68 0.65 0.61
262 32 6 0.0687 1145.30 21.79 0.66 0.65
@@
$M_{w}$ BPR $N_{\mathrm{rings}}$ $D_{COM}\zeta/k_{B}T$ $\tau_{R}/\tau$ $R_{g}/a$ Asphericity Prolateness
40 8 5 0.2396 14.91 4.61 0.64 0.57
80 8 10 0.1685 53.74 7.36 0.67 0.66
168 8 21 0.1122 217.18 12.09 0.66 0.63
344 8 43 0.0745 835.88 19.32 0.65 0.60
520 8 65 0.0590 1753.85 24.89 0.64 0.58
160 16 10 0.1133 172.67 10.84 0.65 0.60
240 24 10 0.0898 341.97 13.57 0.62 0.52
320 32 10 0.0751 611.21 16.60 0.65 0.60
@@ Appendix B Additional Shear Flow Properties
Figure B.1: Polyrotaxane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Polyrotaxanes with 40, 80, and 120 backbone beads ($N_{bb}$) are shown in blue, green, and red, respectively; darker shades indicate higher ring density ($N_{\mathrm{rings}}/N_{bb}$). The rings in the polyrotaxanes are composed of 8 beads, regardless of backbone length.
To calculate tumbling time, we monitor the angle between end-to-end vector of the polymers with the x-axis as it progresses through the four quadrants of a full $360^{\circ}$ rotation cycle and record the time taken for the chains to complete each full cycle (i.e., one complete rotation through all four quadrants) throughout the simulation and use the mean value as the tumbling time. We use the radius of gyration along the flow and gradient directions to calculate the orientation angle $\theta$ to measure the alignment of the chains in the $xy$ plane using the following equation Schroeder et al. (2005).
$\tan(2\theta)=\frac{2R_{g,xy}}{R_{g,xx}-R_{g,yy}}$
According to Fig. B1c, the orientation angle ($\theta$) in polyrotaxanes follows a trend similar to linear polymers, showing that the threaded rings have minimal effect on polymer orientation angle.
@@
Figure B.2: Daisy chain properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, and (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Daisy chains with 8, 16, and 32 beads per segment (BPS) are shown in blue, green, and red, respectively; darker shades indicate more segments ($N_{\mathrm{seg}}$). The rings in all daisy chains are composed of 8 beads.
As shown by the orientation angle in Fig. B2c, daisy chains exhibit little orientation in weak flows and reach their peak orientation angle at $\mathrm{Wi}\approx 7$, compared to $\mathrm{Wi}\approx 3$ for linear polymers, indicating that a stronger flow is needed to align daisy chains. This can be attributed to rearrangements at the mechanical bonds. At high $\mathrm{Wi}$, linear polymers reach a small angle of $\approx 2^{\circ}$, while daisy chains of equal molecular weight maintain a higher angle of $\approx 4.5^{\circ}$.
@@
Figure B.3: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (f) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, and (c) orientation angle ($\theta$). Linear polymers are shown in shades of gray with darker shades indicating more beads. Colors from cold to hot indicate higher number of rings ($N_{\mathrm{rings}}$) with fixed beads per ring ($\mathrm{BPR}=8$).
According to Fig. B3c, at low Wi polymers remain unaligned. At moderate Wi, the orientation angle $\theta$ first rises then falls, with polycatenanes of 8 beads per ring exhibiting a secondary increase previously noted in other systems Dutta and Sing (2024). A terminal decrease, reflecting alignment with very strong flow, could not be verified here because chains broke at high Wi. Having more or smaller rings shifts the primary peak to higher Wi and brings the secondary rise to lower Wi. Having more or larger rings make the orientation angle approach that of a linear polymer.
@@
Figure B.4: Polycatenane properties in shear flow as a function of Weissenberg number (Wi): (a) fractional extension in gradient direction $(\langle Y_{bb}\rangle/L)$, (b) fractional extension in cross direction $(\langle Z_{bb}\rangle/L)$, and (c) orientation angle ($\theta$). Darker shades indicate higher beads per ring (BPR) in polycatenanes with fixed number of rings ($N_{\mathrm{rings}}=10)$.
@@
Figure B.5: Comparison of linear and MIP polymers of similar maximum contour length (L) in extensional (a, b) and shear (c, d) flow. Fractional extension in the flow direction ($\langle X_{bb}\rangle/L$) for extension (a) and shear (c) along with extensional viscosity ($\eta_{e}$) (b) and shear viscosity ($\eta$) are given as a function of Weissenberg number (Wi). The linear polymer (black) has 80 backbone beads. The polyrotaxane (blue) has 80 backbone beads including 2 capping beads and 16 rings. The daisy chain (red) has 10 linear segments of 8 beads each interlocked with 18 8 bead rings and includes 2 capping beads. The polycatenane (green) has 10 interconnected rings of 24 beads each.
Appendix C trajectory snapshots
                  Refer to captionLinear (equilibrium)
                  Refer to captionPolyrotaxane (equilibrium)
                  Refer to captionDaisy Chain (equilibrium)
                  Refer to captionPolycatenane (equilibrium)
                  Refer to captionLinear (extension)
                  Refer to captionPolyrotaxane (extension)
                  Refer to captionDaisy Chain (extension)
                  Refer to captionPolycatenane (extension)
                  Refer to captionLinear (Shear 1)
                  Refer to captionLinear (Shear 2)
                  Refer to captionLinear (Shear 3)
                  Refer to captionLinear (Shear 4)
                  Refer to captionLinear (Shear 5)
                  Refer to captionPolyrotaxane (Shear 1)
                  Refer to captionPolyrotaxane (Shear 2)
                  Refer to captionPolyrotaxane (Shear 3)
                  Refer to captionPolyrotaxane (Shear 4)
                  Refer to captionPolyrotaxane (Shear 5)
                  Refer to captionDaisy Chain (Shear 1)
                  Refer to captionDaisy Chain (Shear 2)
                  Refer to captionDaisy Chain (Shear 3)
                  Refer to captionDaisy Chain (Shear 4)
                  Refer to captionDaisy Chain (Shear 5)
                  Refer to captionPolycatenane (Shear 1)
                  Refer to captionPolycatenane (Shear 2)
                  Refer to captionPolycatenane (Shear 3)
                  Refer to captionPolycatenane (Shear 4)
                  Refer to captionPolycatenane (Shear 5)
Figure C.1: Snapshots of polymer trajectories: equilibrium, extensional flow at Wi=10, and shear flow at Wi=100. Shear flow snapshots show a complete tumbling cycle
Appendix D Probability density distributions of ring positions in polyrotaxanes
Figure D.1: Probability density distributions of ring positions in polyrotaxanes subjected to (a) extensional and (b) shear flow as functions of the Weissenberg number (Wi). The normalized coordinate ranges from 1-1 to +1+1, corresponding to the two ends of the polymer backbone. The graphs at $\mathrm{Wi}=0$ in (a) indicate equilibrium distributions.
Appendix E Power-law exponents
Table E.1: Power-law exponents ($\beta$) of shear viscosity for different polymer architectures
Architecture $M_{w}$ $N_{bb}$ BPS BPR $N_{\mathrm{rings}}$ $N_{\mathrm{seg}}$ $\beta$
Linear 40 - - - - - -0.348 $\pm$ 0.012
Linear 80 - - - - - -0.380 $\pm$ 0.009
Linear 120 - - - - - -0.390 $\pm$ 0.019
Polyrotaxane 56 40 - - 2 - -0.268 $\pm$ 0.028
Polyrotaxane 72 40 - - 4 - -0.198 $\pm$ 0.028
Polyrotaxane 104 40 - - 8 - -0.140 $\pm$ 0.023
Polyrotaxane 112 80 - - 4 - -0.290 $\pm$ 0.019
Polyrotaxane 144 80 - - 8 - -0.238 $\pm$ 0.024
Polyrotaxane 208 80 - - 16 - -0.225 $\pm$ 0.028
Polyrotaxane 168 120 - - 6 - -0.294 $\pm$ 0.018
Polyrotaxane 216 120 - - 12 - -0.283 $\pm$ 0.023
Polyrotaxane 312 120 - - 24 - -0.239 $\pm$ 0.018
Daisy chain 30 - 8 - - 2 -0.152 $\pm$ 0.042
Daisy chain 74 - 8 - - 4 -0.207 $\pm$ 0.018
Daisy chain 118 - 8 - - 6 -0.219 $\pm$ 0.009
Daisy chain 206 - 8 - - 10 -0.212 $\pm$ 0.013
Daisy chain 46 - 16 - - 2 -0.245 $\pm$ 0.020
Daisy chain 106 - 16 - - 4 -0.252 $\pm$ 0.011
Daisy chain 166 - 16 - - 6 -0.267 $\pm$ 0.030
Daisy chain 226 - 16 - - 8 -0.275 $\pm$ 0.017
Daisy chain 78 - 32 - - 2 -0.326 $\pm$ 0.021
Daisy chain 170 - 32 - - 4 -0.317 $\pm$ 0.017
Daisy chain 262 - 32 - - 6 -0.325 $\pm$ 0.016
Polycatenane 40 - - 8 5 - -0.158 $\pm$ 0.024
Polycatenane 80 - - 8 10 - -0.154 $\pm$ 0.016
Polycatenane 168 - - 8 21 - -0.163 $\pm$ 0.046
Polycatenane 344 - - 8 43 - -0.172 $\pm$ 0.021
Polycatenane 520 - - 8 65 - -0.203 $\pm$ 0.039
Polycatenane 160 - - 16 10 - -0.238 $\pm$ 0.019
Polycatenane 240 - - 24 10 - -0.298 $\pm$ 0.009
Polycatenane 320 - - 32 10 - -0.306 $\pm$ 0.014
Table E.2: Power-law exponents ($\beta$) of $\psi_{1}$ for different polymer architectures
Architecture $M_{w}$ $N_{bb}$ BPS BPR $N_{\mathrm{rings}}$ $N_{\mathrm{seg}}$ $\beta$
Linear 40 - - - - - -0.976 $\pm$ 0.013
Linear 80 - - - - - -1.239 $\pm$ 0.018
Linear 120 - - - - - -1.317 $\pm$ 0.009
Polyrotaxane 56 40 - - 2 - -1.052 $\pm$ 0.016
Polyrotaxane 72 40 - - 4 - -0.959 $\pm$ 0.016
Polyrotaxane 104 40 - - 8 - -0.912 $\pm$ 0.013
Polyrotaxane 112 80 - - 4 - -1.178 $\pm$ 0.023
Polyrotaxane 144 80 - - 8 - -1.125 $\pm$ 0.023
Polyrotaxane 208 80 - - 16 - -1.183 $\pm$ 0.025
Polyrotaxane 168 120 - - 6 - -1.205 $\pm$ 0.030
Polyrotaxane 216 120 - - 12 - -1.283 $\pm$ 0.025
Polyrotaxane 312 120 - - 24 - -1.240 $\pm$ 0.023
Daisy chain 30 - 8 - - 2 -0.947 $\pm$ 0.032
Daisy chain 74 - 8 - - 4 -1.105 $\pm$ 0.018
Daisy chain 118 - 8 - - 6 -1.199 $\pm$ 0.011
Daisy chain 206 - 8 - - 10 -1.255 $\pm$ 0.015
Daisy chain 46 - 16 - - 2 -1.225 $\pm$ 0.017
Daisy chain 106 - 16 - - 4 -1.320 $\pm$ 0.014
Daisy chain 226 - 16 - - 8 -1.423 $\pm$ 0.019
Daisy chain 78 - 32 - - 2 -1.349 $\pm$ 0.033
Daisy chain 170 - 32 - - 4 -1.415 $\pm$ 0.023
Daisy chain 262 - 32 - - 6 -1.497 $\pm$ 0.033
Polycatenane 40 - - 8 5 - -0.918 $\pm$ 0.024
Polycatenane 80 - - 8 10 - -1.067 $\pm$ 0.023
Polycatenane 168 - - 8 21 - -1.342 $\pm$ 0.037
Polycatenane 344 - - 8 43 - -1.499 $\pm$ 0.025
Polycatenane 520 - - 8 65 - -1.602 $\pm$ 0.027
Polycatenane 160 - - 16 10 - -1.064 $\pm$ 0.027
Polycatenane 240 - - 24 10 - -1.172 $\pm$ 0.020
Polycatenane 320 - - 32 10 - -1.197 $\pm$ 0.027